2X + 4X +10 - 3 = 28
Answer by Nieleya(22)

2X + 4X +10 - 3 = 28
6x + 7 = 28
6x = 28 - 7
6x = 21
x = 21/6
x = 3.5
Proof
2 (3.5) + 4(3.5) + 10 - 3
7 + 14 + 10 - 3
= 28